1. Between z = 0 and z = 1.36
1
Expert's answer
2022-03-01T09:17:49-0500


The area between Z = 0 and Z = 1.36 is computed using the formula:

P(0<Z<1.36) = P(Z<1.36) - P(Z<0)

From the standard normal table,

P(Z<1.36) = 0.9131

Also,

P(Z<0) = 0.5

Therefore,

P(0<Z<1.36) = P(Z<1.36) - P(Z<0) = (0.9131 - 0.5O) = 0.4131


Answer: 0.4131